# mO shizhe[#1]

### v1

@@ 5.1 Hallucination Detection Algorithms
Hallucination Detection Prompt: Hallucination detection prompts are manually crafted to instruct LLMs (GPT-4-turbo and GPT-3.5-turbo) in assessing whether a given reference-response pair contains hallucinated content and to identify the corresponding hallucinated spans in the response. For detailed information about these prompts, please refer to Appendix C.
Methods Question Answering Data-to-text Summarization Overall
@@
Methods Question Answering Data-to-text Summarization Overall
@@
SelfCheckGPT (Manakul et al., 2023): SelfCheckGPT employs a zero-resource, sampling-based method to fact-check the responses of black-box models. When processing each response in RAGTruth, 5 responses from other models serve as references, and GPT-4-turbo/GPT-3.5-turbo are queried to verify consistency. By scanning each sentence in a response, the hallucination position can be determined as well.
@@
LLM Finetuning: Llama-2-13B has been fine-tuned using the training set from RAGTruth. The model takes the context-response pair with proper instructions as the input and treats the hallucinate span as the targeted generation output. We employed LoRA Hu et al. (2021) fine-tuning with an initial learning rate of 3e-4, and limiting the training to 3 epochs, all conducted on 4 A100 GPUs.
@@ 5.3 Evaluation Metrics
We calculate the overlap between the generated span and human-labeled span and report the precision, recall, and f1 score at the char-level.
@@ 6.1 Response-level Detection
The results in Table 5 reveal that hallucination detection remains a significant challenge in the context of RAG for all existing detection methods. Even when reference information is available, the responses generated may still include hallucinations, which current LLMs cannot reliably identify. The most advanced LLM, GPT-4-turbo, achieves only an average F1 score of 68.3%. In comparison, the widely-used GPT-3.5-turbo often fails to discern hallucinated responses, often identifying most examples as non-hallucinated, highlighting the difficulty in detecting naturally occurring hallucinations. For another notable baseline, SelfCheckGPT also shows unsatisfactory performance in this regard, achieving an average F1 score of 60.5% with GPT-4-turbo and 36.6% with GPT-3.5-turbo.
By utilizing our human-labeled high-quality training set, a fine-tuned Llama-2-13B can achieve the best performance with an average 80.7% f1 score. This shows the effectiveness of our data in improving the model’s detection ability on hallucination.
@@ 6.2 Span-level Detection
RAGTruth, as a hallucination corpus with fine-grained span labels, enables us to present experimental results for span-level detection, serving as a baseline for future research. As shown in Table 6, the overall performance of the current detection method is sub-optimal, highlighting the challenges in span-level detection. Even the advanced GPT-4-turbo tends to incorrectly classify many non-hallucinated contents with a low precision of 24%. While our fine-tuned model shows improved capability in identifying hallucinated spans by achieving an averaged f1 score of 48.4%, it still falls short of perfect detection, emphasizing the inherent difficulties of this task.
Group Treatment Valid Response Num Hallucination Rate
1 Sample From Llama-2-7B-chat 450 51.8
Mistral-7B-Instruct 450 57.6
Sampling Strategy Random selection 450 55.1(-)
Least detected spans 450 43.1($\downarrow$21.8%)
No detected spans ${}^{\dagger}$ 23.9($\downarrow$56.6%)
2 Sample From GPT-3.5-Turbo-0613 450 10.9
GPT-4-0613 450 9.3
Sampling Strategy Random selection 450 10.4(-)
Least detected spans 450 5.3($\downarrow$41.0%)
No detected spans ${}^{\dagger}$ 4.9($\downarrow$52.9%)
@@ 6.3 Hallucination Suppression
We tested the effectiveness of hallucination suppression using our finetuned hallucination detection model. For the 450 instances in the test set, we employed two strategies to select a final output from two responses generated by two different models with similar hallucination density: one strategy is based on choosing the response with fewer predicted hallucination spans, and the other is based on the absence of detected hallucination span. When the number of hallucination spans detected in both candidate responses is the same, one will be chosen at random. Due to limited response candidates, not all instances have a response that conforms to the second strategy. In practical scenarios, this issue can be addressed by increasing the number of candidate responses. We employed random selection as a simple baseline for comparison.
The results shown in Table 7 indicate that with the help of the hallucination detector, both strategies can significantly reduce hallucination rate. For the relatively small Llama-2-7B-chat and Mistral-7B-Instruct models, compared to random selection, the first strategy reduced the hallucination rate by 21.8%, while the second strategy achieved a reduction of 56.6%. Even for models with a low hallucination rate, specifically GPT-3.5-Turbo and GPT-4, employing the finetuned hallucination detector for sampling can still further reduce the rate of hallucinations. The two strategies yielded a reduction in hallucination rates of 41% and 52.9%, respectively. These results demonstrate the potential of an efficient hallucination detection model in developing trustworthy RAG LLMs.
@@ 7 Conclusion
In this paper, we introduce RAGTruth, a large-scale corpus of naturally generated hallucinations, featuring detailed word-level annotations tailored for Retrieval-augmented Generation (RAG) scenarios. Our work includes an in-depth analysis of the interplay between hallucinations and various factors, such as task types, models being used, and contextual settings.
Additionally, we conduct empirical benchmarks of several hallucination detection approaches using our corpus. We show that fine-tuning Llama with RAGTruth leads to competitive performance in hallucination detection. This implies that by using a high-quality dataset such as RAGTruth, it is possible to develop specialized hallucination detection models that are highly effective when compared to prompt-based methods using general models such as GPT-4. The dataset paves the way for future research in this direction.
Simultaneously, our findings reveal that identifying hallucinations in RAG contexts, particularly at the span level, remains a formidable challenge, with current methods still falling short of reliable detection. We hope that RAGTruth, with its comprehensive span annotations, can assist the development of hallucination detection techniques for retrieval augmented generation.
@@ 8 Acknowledgement
We appreciate the valuable feedback and assistance from Shizhe Diao. We thank Dongfang Li for her support in illustration.
@@ Appendix A Hallucination Examples
Original content: in a restaurant review where a customer briefly mentions a "happy hour" without specifying the timing
Generative content: assumes it to be between 4-7 PM.
@@ Appendix B Response Generation Prompts
Summarization
Summarize the following news within {word_num} words:
{news}
output:
@@
Table 9: Prompts for generating responses for the three types of tasks. word_num is min(200, word_num_of_news//4)
Appendix C Hallucination Detection Prompts
@@
Article: {context}
Summary: {sentence}
Please analyze the provided summary and indicate whether there are any hallucinated (invented or incorrect) parts. If there are, please specify the exact span that is hallucinated.
Please provide a response in the following format:
{"hallucination_detected": "yes/no",
"hallucinated_span": "If yes, the exact span of every hallucinated text part in list format; if no, leave this empty."}
@@
Passages: {passage}
Question: {question}
Answer: {answer}
Please analyze the provided Passages, Question, and Answer, and indicate whether there are any hallucinated (invented or incorrect) parts in the answer. If there are, please specify the exact span that is hallucinated.
Please provide a response in the following format:
{"hallucination_detected": "yes/no",
"hallucinated_span": "If yes, the exact span of every hallucinated text part in list format; if no, leave this empty."}
Data-to-text
Information: {information}
Overview: {response}
Please analyze the provided business information and its overview, and indicate whether there are any hallucinated (invented or incorrect) parts in the answer. If there are, please specify the exact span that is hallucinated.
Please provide a response in the following format:
{"hallucination_detected": "yes/no",
"hallucinated_span": "If yes, the exact span of every hallucinated text part in list format; if no, leave this empty."}
Table 10: Prompts for detecting hallucination for the three types of tasks.

### v2

@@ 5.1 Hallucination Detection Algorithms
Hallucination Detection Prompt: Hallucination detection prompts are manually crafted to instruct LLMs (GPT-4-turbo and GPT-3.5-turbo) in assessing whether a given reference-response pair contains hallucinated content and to identify the corresponding hallucinated spans in the response. For detailed information about these prompts, please refer to Appendix D.
Methods Question Answering Data-to-text Writing Summarization Overall
@@
Methods Question Answering Data-to-text Writing Summarization Overall
@@
SelfCheckGPT (Manakul et al., 2023): SelfCheckGPT employs a zero-resource, sampling-based method to fact-check the responses of black-box models. When processing each response in RAGTruth, 3 extra responses from the same model were sampled and served as references, and GPT-3.5-turbo was used to verify consistency. We detected hallucinations sentence-by-sentence within a response, and then aggregated these results to provide a response-level detection outcome.
@@
LLM Finetuning: Llama-2-13B has been fine-tuned using the training set from RAGTruth. The model takes the context-response pair with proper instructions as the input and treats the hallucinate span as the targeted generation output. We employed full training with an initial learning rate of 2e-5, and limiting the training to 1 epochs, all conducted on 4 A100 GPUs.
@@ 5.3 Evaluation Metrics
It is a more challenging and significant task to identify the locations of hallucinations within the response than only determining whether a response contains hallucinations. We assess hallucination detection at both the response and span levels.
@@
We calculate the overlap between the detected span and human-labeled span and report the precision, recall, and f1 score at the char-level.
@@ 6.1 Response-level Detection
The results in Table 5 reveal that hallucination detection remains a significant challenge in the context of RAG for all existing detection methods. Even when reference information is available, the responses generated may still include hallucinations, which current LLMs cannot reliably identify. The most advanced LLM, GPT-4-turbo, achieves only an average F1 score of 63.4%. For another notable baseline, SelfCheckGPT also shows unsatisfactory performance in this regard, achieving an average F1 score of 58.8% with GPT-3.5-turbo.
By utilizing our high-quality training set, a fine-tuned Llama-2-13B can achieve the best performance with an average 78.7% f1 score. This shows the effectiveness of our data in improving the model’s hallucination detection ability.
@@ 6.2 Span-level Detection
RAGTruth, as a hallucination corpus with fine-grained span labels, enables us to present experimental results for span-level detection, serving as a baseline for future research. As shown in Table 6, the overall performance of the current detection method is sub-optimal, highlighting the challenges in span-level detection. Even the advanced GPT-4-turbo tends to incorrectly classify many non-hallucinated contents with a low precision of 18.4%. While our fine-tuned model shows improved capability in identifying hallucinated spans by achieving an averaged f1 score of 52.7%, it still falls short of perfect detection, emphasizing the inherent difficulties of this task.
Group Selection Strategy Valid Response Num Hallucination Rate
Llama-2-7B-chat (51.8) Mistral-7B-Instruct (57.6) Random 450 52.4(-)
Select the response with fewer detected hallucination spans 450 41.1($\downarrow$21.6%)
Select the response with no detected hallucination spans 328 19.3($\downarrow$63.2%)
GPT-3.5-Turbo-0613 (10.9) GPT-4-0613 (9.3) Random 450 9.8(-)
Select the response with fewer detected hallucination spans 450 5.6($\downarrow$42.9%)
Select the response with no detected hallucination spans 448 4.8($\downarrow$51.0%)
@@ 6.3 Hallucination Suppression
We tested the effectiveness of hallucination suppression using our finetuned hallucination detection model. For the 450 instances in the test set, we employed two strategies to select a final output from two responses generated by two different models with similar hallucination densities. The first strategy involved selecting the response with fewer predicted hallucination spans. The second strategy, more stringent, mandated that the selected response have no detected hallucination spans. When the number of hallucination spans detected in both candidate responses is the same, one will be chosen at random. Due to limited response candidates, not all instances have a response that conforms to the second strategy. In practical scenarios, this issue can be addressed by increasing the number of candidate responses. We employed random selection as a simple baseline for comparison.
The results shown in Table 7 indicate that with the help of the hallucination detector, both strategies can significantly reduce the hallucination rate. For the relatively small Llama-2-7B-chat and Mistral-7B-Instruct models, compared to random selection, the first strategy reduced the hallucination rate by 21.6%, while the second strategy achieved a reduction of 63.2%. Even for models with a low hallucination rate, specifically GPT-3.5-Turbo and GPT-4, employing the finetuned hallucination detector for sampling can still further reduce the rate of hallucinations. The two strategies yielded a reduction in hallucination rates of 42.9% and 51.0%, respectively. These results demonstrate the potential of an efficient hallucination detection model in developing trustworthy RAG LLMs.
@@ 7 Conclusion
In this paper, we introduce RAGTruth, a large-scale corpus of naturally generated hallucinations, featuring detailed word-level annotations tailored for RAG scenarios. Our work includes an in-depth analysis of the interplay between hallucinations and various factors, such as task types, models being used, and contextual settings.
Additionally, we conduct empirical benchmarks of several hallucination detection approaches using our corpus. We show that fine-tuning Llama with RAGTruth leads to competitive performance. This implies that by using a high-quality dataset such as RAGTruth, it is possible to develop specialized hallucination detection models that are highly effective when compared to prompt-based methods using general models such as GPT-4.
Simultaneously, our findings reveal that identifying hallucinations in RAG contexts, particularly at the span level, remains a formidable challenge, with current methods still falling short of reliable detection. We hope that RAGTruth, can assist the development of hallucination detection techniques for retrieval augmented generation.
8 Limitations
The study of hallucination in large language models is a rapidly advancing field, characterized by the continuous evolution of application scenarios, sources of hallucination, and techniques for detecting and preventing them. While our work represents the first attempt to benchmark hallucination in the RAG setting, there may be situations not addressed by this research that are nonetheless significant for certain practical applications.
9 Ethical considerations
This work is in full compliance with the Ethics Policy of the ACL. We acknowledge that responses generated by LLMs in this study may contain inaccuracies. Aside from this, to the best of our knowledge, there are no additional ethical issues associated with this paper.
@@ 10 Acknowledgement
We appreciate the valuable feedback and assistance from Shizhe Diao. We thank Doris Li for her support in creating the illustrations for this research.
@@ Appendix A Hallucination Examples
Example 2
Original content: "RestaurantsReservations": false
Generative content: The restaurant offers outdoor seating and takes reservations
@@
Example 2
Original content: It was quite a hassle to find a table since this place is crowded all the time, but well, nothing can stand between me and that absolutely delightful lobster roll
Generative content: The restaurant is located in a crowded area, but the delicious lobster rolls make it worth the hassle to find a table.
@@
Example 2
Original content: “RestaurantsTakeOut”: true
Generative content: Customers can also enjoy takeout and delivery options.
@@
Example 1
Original content: Amazing happy hour and great people.
Generative content: One of the standout features of Killer Shrimp is their happy hour, which is available daily from 4 pm to 7 pm.
Example 2
Original content: Place mashed potatoes in a medium bowl. Stir in onions, cheese, salt, pepper, and bacon.
Generative content: Place your mashed potatoes in a medium bowl. Stir in cooked onions, cheese, salt, pepper, and bacon.
@@ Appendix B Response Generation Prompts
Summarize the following news within {word_num} words:
{news}
output:
Table 9: Prompts for generating responses for the three types of tasks. word_num is min(200, word_num_of_news//4). The word count requirement is only to control the length of the generated summarization, it will not serve as the basis for hallucination annotation.
Appendix C Annotation Details
Figure 5: Annotation interface. For privacy reasons, we have masked the full names of the annotators in the screenshot.
Task Model # Hallucination Span implicit_true due_to_null
# Span % Span # Span % Span
Question Answering GPT-3.5-turbo-0613 89 33 0.371
GPT-4-0613 51 15 0.294
Llama-2-7B-chat 1010 251 0.249
Llama-2-13B-chat 654 215 0.329
Llama-2-70B-chat 529 168 0.318
Mistral-7B-Instruct 594 164 0.276
Data-to-text Writing GPT-3.5-turbo-0613 384 52 0.135 69 0.180
GPT-4-0613 354 24 0.068 209 0.590
Llama-2-7B-chat 1775 195 0.110 230 0.130
Llama-2-13B-chat 2803 260 0.09 439 0.157
Llama-2-70B-chat 1834 274 0.149 272 0.148
Mistral-7B-Instruct 2140 102 0.048 423 0.198
Summarization GPT-3.5-turbo-0613 60 14 0.233
GPT-4-0613 80 10 0.125
Llama-2-7B-chat 517 44 0.085
Llama-2-13B-chat 342 28 0.082
Llama-2-70B-chat 245 27 0.110
Mistral-7B-Instruct 828 52 0.063
Overall 14289 1928 0.135 1642 0.115
Table 10: Detailed statistical information for the labels implicit_true and due_to_null. The majority of implicit truths appear in two types of tasks: question answering and data-to-text writing. About 17.7% hallucination spans in the data-to-text writing tasks are related to null values in the JSON data.
Appendix D Hallucination Detection Prompts
Summarization
Below is the original news:
{article}
Below is a summary of the news:
{summary}
Your task is to determine whether the summary contains either or both of the following two types of hallucinations:
1. conflict: instances where the summary presents direct contraction or opposition to the original news;
2. baseless info: instances where the generated summary includes information which is not substantiated by or inferred from the original news.
Then, compile the labeled hallucinated spans into a JSON dict, with a key "hallucination list" and its value is a list of hallucinated spans. If there exist potential hallucinations, the output should be in the following JSON format: {"hallucination list": [hallucination span1, hallucination span2, …]}. Otherwise, leave the value as a empty list as following: {"hallucination list": []}.
Output:
@@ Appendix D Hallucination Detection Prompts
Below is a question:
{question}
Below are related passages:
{passages}
Below is an answer:
{answer}
Your task is to determine whether the answer contains either or both of the following two types of hallucinations:
1. conflict: instances where the answer presents direct contraction or opposition to the passages;
2. baseless info: instances where the answer includes information which is not substantiated by or inferred from the passages.
Then, compile the labeled hallucinated spans into a JSON dict, with a key "hallucination list" and its value is a list of hallucinated spans. If there exist potential hallucinations, the output should be in the following JSON format: {"hallucination list": [hallucination span1, hallucination span2, …]}. Otherwise, leave the value as a empty list as following: {"hallucination list": []}.
Output:
Data-to-text Writing
Below is a structured data in the JSON format:
{business info}
Below is an overview article written in accordance with the structured data:
{overview}
Your task is to determine whether the overview contains either or both of the following two types of hallucinations:
1. conflict: instances where the overview presents direct contraction or opposition to the structured data;
2. baseless info: instances where the generated overview includes information which is not substantiated by or inferred from the structured data.
In JSON, "null" or "None" represents an unknown value rather than a negation.
Then, compile the labeled hallucinated spans into a JSON dict, with a key "hallucination list" and its value is a list of hallucinated spans. If there exist potential hallucinations, the output should be in the following JSON format: {"hallucination list": [hallucination span1, hallucination span2, …]}. Otherwise, leave the value as a empty list as following: {"hallucination list": []}.
Output:
Table 11: Prompts for detecting hallucination for the three types of tasks. In the prompt for data-to-text writing, we clarified that null or None in JSON should be treated as unknown rather than a negation.